Write a gradient of a straight line the pass through the points (2,3) and (-1,2)
1
Expert's answer
2020-10-02T15:58:37-0400
"grad=\dfrac{y_2-y_1}{x_2-x_1}=\dfrac{2-3}{-1-2}=\dfrac{1}{3}"